20230204 UT Second Half ECD

Night center: 00:33 MST

Morning 12-deg twilight: 06:17 MST

Sunrise: 07:09 MST

Last night: No DX AO work. SX-ESM testing with DX-SL imaging. Vignetting in AGw1 explored (IT 8495)

Night summary: The seeing was good (0.6-0.8") and binocular bin 1 ESM tests were done. This was the first DX ESM observation. Then some DX-SOUL COMM (bin 4) was attempted, but clouds moved in making that difficult. Reconfigured to the LBCs for on-sky tests of channel-dropping. The temperature swings made collimation very difficult, but eventually some channel-dropping tests were made although not the full test to the minimum channel configuration. There were issues with copointing limits during the LUCI observations which were resolved by restarting PCS. On the first LBC preset, mirror limits were hit due to the large mode 3 pointing offsets (Y ~ 2mm). Resolution involved restarting PCS (though this did not help), reauthorizing the LBCs and dropping back to the default IE/CA.

Tasks to be done:

Google doc spreadsheet with outline of plan

Repoint (0.5hours (more if closed dome or cloudy may be used))

  • Test the repoint command in the AOS
  • Test repoint command through scripts
  • Repoint moves the telescope to place the Ref Star on the hotspot (currently does not work)
    • Verify amplitude is correct
    • Verify Direction is correct
    • Verify angle is correct

Bin1 ESM check (1hr)

  • Testing using the ESM_V57.xml OT, testing regular ESM optimization and data quality
  • LUCI1 ESM and LUCI2 SL, LUCI2 is ready for ESM so may run fully binocular.
  • Targets M61, M67
  • Seeing <2"

SESM Bin1 reconstructor 60 modes vs 150modes (2hr)

  • LUCI1-SESM & LUCI2-SL or for fun LUCI1_SESM and LUCI2 ESM
  • ESM_V57.xml M92 or M67
  • Compare 150 mode reconstructor with 50 mode reconstruction on full N375 field on sky

TMS testing (3hr)

  • After the power outage, Heejoo found the same channels as before, but some were weaker. Pat worked on alignment and was able to bring back two channels (5 & 20) that were noisy and had often been filtered out during the recent LBC/TMS use on Jan 26, 27 UT. Results summarized by this figure which Heejoo provided:
  • TMSalignment_20230203.png:
    TMSalignment_20230203.png
  • Channel-dropping on-sky with TMS updates mid-exposure:
    • selectively drop and bring back RR pairs and monitor the effect of the different configurations (see attached TMSchannelschema.png)
    • tms_loop.py -o (-o is needed to ignore the shutter state and send corrections during exposures).
    • single filter, no dithering, series of 30-60 sec exposures, fields with fairly high star density at low & high elevation.
  • [DEFER] Thermal aberrations
    • need to analyze current data first, but a proposed future test would be to collect extra-focal pupil images while driving mirror temperatures. Requires ~6 hrs and seeing < 1.4". Clouds OK
  • TMS channel schema: collimator-retroreflector pairs. Note that in the image below:
    • collimator #29 corresponds to TMS channel #19; and
    • collimator #30 corresponds to TMS channel #21
    • TMSchannelschema.png
  • NGC4147 - elev= 75-60 deg at 10:45-12:00 UT - collimate and start here (high-mid elevation field)
  • 10h85d - elev = 38-35 deg at 10:45-12:00 UT - low elevation field
  • NGC5904 - elev= 41-58 deg at 11-13:00 UT
  • NGC5466 - above 80 deg elev between 11:50-13:15
  • NGC5272 - above 80 deg elev between 11:25-12:45

Bin4 red targets

  • Clear, thing cirrus with seeing <2" * ESM_V57.xml M92 or M67

Check edges of LUCI-AGw patrol fields (3.5 hrs?)

  • Some investigation done last night
  • SL: both LUCI1 and LUCI2
  • poor weather
  • target prepared: NGC3190 @ RA=10, DEC=+21

Night Log

Link to google doc log:

https://docs.google.com/document/d/1pcCuMW_s_QRnIM16h7E8CnCsqlsTv_3t9Qgiu7T03H4/edit?usp=sharing

-- OlgaKuhn - 03 Feb 2023
I Attachment Action Size Date Who Comment
TMSalignment_20230203.pngpng TMSalignment_20230203.png manage 2 MB 03 Feb 2023 - 22:32 OlgaKuhn  
TMSchannelschema.pngpng TMSchannelschema.png manage 284 K 03 Feb 2023 - 20:48 OlgaKuhn TMS channel schema: collimator-retroreflector pairs
Topic revision: r8 - 05 Feb 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback